Find the point(s) on the graph of x^2+y^2=45 at which the slope of the tangent is -1/2.
Answer by Alan3354(69443)

Find the point(s) on the graph of x^2+y^2=45 at which the slope of the tangent is -1/2.
-------------
slope m = -x/y
---
-x/y = -1/2
y = 2x
=============
x^2+y^2=45
x^2+(2x)^2=45
5x^2 = 45
x^2 = 9
x = +3, y = 6 --> (3,6)
x = -3, y = 6 --> (-3,6)